# Too Many Frames, Not All Useful: Efficient Strategies for Long-Form Video QA

### v5

@@ Abstract
Long-form videos that span across wide temporal intervals are highly information redundant and contain multiple distinct events or entities that are often loosely related. Therefore, when performing long-form video question answering (LVQA), all information necessary to generate a correct response can often be contained within a small subset of frames. Recent literature explore use of large language models (LLMs) in LVQA benchmarks, achieving exceptional performance, while relying on vision language models (VLMs) to convert all visual content within videos into natural language. Such VLMs often independently caption a large number of frames uniformly sampled from long videos, which is not efficient and can mostly be redundant. Questioning these decision choices, we explore optimal strategies for key-frame selection that can significantly reduce these redundancies, namely Hierarchical Keyframe Selector. Our proposed framework, LVNet, achieves state-of-the-art performance at a comparable caption scale across three benchmark LVQA datasets: EgoSchema, NExT-QA, and IntentQA, while also demonstrating a strong performance on videos up to an hour long in VideoMME. The code can be found at https://github.com/jongwoopark7978/LVNet
@@ 1 Introduction
Video understanding is a long-standing vision problem (Aggarwal and Ryoo, 2011) with numerous real-world applications. It has been traditionally studied even before the era of differentiable representation learning, with hierarchical approaches focusing on longer videos (Allen and Ferguson, 1994; Ivanov and Bobick, 2000; Shi et al., 2004; Hongeng et al., 2004; Ryoo and Aggarwal, 2006). Today, video understanding research involving the language modality is particularly popular, with tasks such as video question answering that involve generating human-style conversations in response to questions regarding videos (Tapaswi et al., 2016; Zeng et al., 2017; Xu et al., 2017).
Figure 1: LVNet achieves state-of-the-art performance on EgoSchema subset while utilizing only a fraction of captioned frames. In particular, LVNet obtains its highest accuracy of 68.2% with 12 captions (VLM calls), outperforming VideoAgent and VideoTree, models using a similar-scale captions, by +8% and +5.7% (more details in Section 4.3).
Recent popularity of vision-language models (VLMs), particularly approaches connecting large language models (LLMs) to vision architectures (Liu et al., 2023; Li et al., 2023b; Dai et al., 2023), has resulted in significant improvements across visual question answering (VQA) tasks. These models demonstrate exceptional performance within the image domain, and their video variants (Yu et al., 2023; Papalampidi et al., 2023; Maaz et al., 2023) perform similarly on shorter videos, yet demonstrate limited performance on long-form video benchmarks (Mangalam et al., 2023; Kahatapitiya et al., 2024; Rawal et al., 2024). This can be attributed to the nature of long-form video benchmarks, which require both temporal sequence awareness and causal reasoning. An alternate line of works (Zhang et al., 2023; Wang et al., 2023; Kahatapitiya et al., 2024; Wang et al., 2024b) adapt LLMs that contain strong reasoning abilities for this task, using image VLMs to generate per-frame natural language descriptions, followed by video question answering purely within the language domain. However, these methods employ expensive VLMs to caption a large number of uniformly sampled frames. Such a design choice leading to high compute expense, is questioned in (Buch et al., 2022; Ranasinghe et al., 2024; Wang et al., 2024b), and is the key motivation for our exploration of key frame selection, i.e. identifying a minimal set of frames most useful for correctly answering a given video-question pair.
@@
Therein, we propose LVNet, a framework containing a novel Hierarchical Keyframe Selector (HKS) that performs efficient key-frame selection followed by VLM and LLM for caption and answer generation as illustrated in Fig. 2. Aligned with prior work (Zhang et al., 2023; Wang et al., 2024d, b), the per-frame captions are processed with a powerful LLM to generate correct answers for a given video-question pair. As shown in Fig. 1, LVNet achieves strong performance using a small set of keyframes from the HKS. The scope of this work focuses on optimizing the prior two stages. We summarize our key contributions as follows:
1. Hierarchical Keyframe Selector (HKS): The proposed HKS consists of three submodules for efficient keyframe selection.
@@
    Temporal Scene Clustering (TSC)
    • Performs non-uniform frame sampling by clustering visually similar frames.
    • Reduces redundancy in long videos while capturing key scenes.
    • A lightweight module for efficient filtering of dense frames.
@@
    Coarse Keyframe Detector (CKD)
    • Generates keywords representing atomic activities using the given query and an LLM.
    • Assigns confidence scores to frames based on keyword relevance.
    • Samples high-confidence frames for improved interpretability over visual-only selection.
@@
    Fine Keyframe Detector (FKD)
    • Refines frame selection by combining multiple frames using visual templating and a VLM.
    • Enables higher-level reasoning and natural language-based selection.
    • Achieves better accuracy than CKD’s keyword-based selection.
2. Zero-Shot Long-Form Video Understanding: Our framework operates zero-shot without requiring video-level training. This makes it highly efficient for long-form video understanding.
@@
Proposed LVNet achieves state-of-the-art results compared to models utilizing similar number of captions on three long-form video question answering benchmarks— EgoSchema, NExT-QA, and IntentQA(Sec. 4.2). This demonstrates strong performance and generality of our approach.
@@ 2 Related Work
Video Question Answering: Visual question answering (VQA) involves generating open-ended textual content conditioned on an image and natural language query (Agrawal et al., 2015). Its video variant, Video-VQA (Yu et al., 2019a) replaces images with videos. Multiple early datasets focus on querying objects or events based on referential and spatial relations (Xu et al., 2017; Zeng et al., 2017; Yu et al., 2019a). Later tasks require explicit temporal understanding of sequential events (Lei et al., 2018, 2020; Yu et al., 2019b). More recent datasets focus on longer videos containing multiple actions and scenes spread over wide time intervals (termed long-form videos) (Xiao et al., 2021; Li et al., 2022). Referred to as long-form video question answering (LVQA), these benchmarks are constructed to specifically test strong causal and temporal reasoning (Xiao et al., 2021) over long temporal windows (Mangalam et al., 2023). Some works tackling such video VQA tasks leverage graph networks to model cross object / event relations (Hosseini et al., 2022; Xiao et al., 2022a, b). A more recent line of works integrate LLMs to tackle this task (Zhang et al., 2023; Wang et al., 2023; Kahatapitiya et al., 2024; Wang et al., 2024b; Ranasinghe et al., 2024; Wang et al., 2024d; Fan et al., 2024) utilizing the strong reasoning skills of LLMs. A common aspect is the use of a vision language model (VLM) to convert frame level visual information into natural language. This in turn is input to the LLM which makes a final prediction.
@@
Table 1: LVNet exhibits unique features compared to VideoAgent (VA) Fan et al. (2024), Traveler (Tr.) Shang et al. (2024) and VideoTree (VT) Wang et al. (2024e). See Appendix A.5 for details.
Frame Selection in Videos: The task of frame selection in videos has been long explored in video (Davis and Bobick, 1997; Zhao et al., 2017) with more recent works focused directly on long-form video question answering (Buch et al., 2022; Wang et al., 2024e; Fan et al., 2024). Most similar to our work is (Wang et al., 2024b) which employs an LLM based strategy for video frame selection. However, our LVNet differs with several unique features as summarized in Table 1.
@@ 3 Method
Our proposed LVNet comprises of three components: a Hierarchical Keyframe Selector (HKS), a Vision Language Model (VLM), and a Large Language Model (LLM) as illustrated in Figure 2. The HKS, an efficient, hierarchical keyframe selector, is the core contribution of our work. First, the model processes 900 uniformly sampled frames and clusters them into distinct scenes Next, it extracts keywords from a given natural language query via LLM and selects the frames most relevant to those keywords. Finally, the selected frames are described in natural language by a more powerful and computationally intensive VLM. Finally, an LLM processes the language descriptions of the selected frames to answer a given query.
@@ 3.1 Background
Recent approaches utilizing LLMs for long video question answering (LVQA) (Zhang et al., 2023; Wang et al., 2023; Kahatapitiya et al., 2024; Ranasinghe et al., 2024; Wang et al., 2024b) can be viewed as a composition of three sequential stages: a) frame selection, b) VLM based frame captioning, and c) LLM based answer generation. Note that the complexity of each stage varies across methods given their focus on different aspects of the LVQA task (e.g. frame selection in some is simply uniform sampling). In our work, we also follow this structure, but we focus on improving the frame selection stage. Under such a framework, our proposed HKS can serve as plug-in modules to replace the frame selection stage and the later two stages are similar to these prior works.
@@ 3.2 Architecture
Our LVNet processes a given video-query ($\mathbf{x},\mathbf{q}$) pair to output a response, $\mathbf{\hat{r}}$. The HKS module initially processes this video-query pair, selects $T^{\prime}$ keyframes, and outputs a deterministically sub-sampled video $\mathbf{x^{\prime}}\in\mathbb{R}^{T^{\prime}\times C\times H\times W}$. Each of these $T^{\prime}$ frames is then passed through the captioning stage of our VLM to generate a set of natural language descriptions, $D=\{d_{1},d_{2},...d_{T^{\prime}}\}$ where $d_{i}$ describes the frame $\mathbf{x^{\prime}}[i]$. Finally, the LLM processes all descriptions $D$ and the query $\mathbf{q}$ to generate response $\mathbf{\hat{r}}$. We illustrate this overall architecture in Figure 2.
@@ 3.3 Hierarchical Keyframe Selector
We now describe our proposed Hierarchical Keyframe Selector (HKS) module. As illustrated in Figure 2, our proposed HKS comprises of three sequential submodules, each reducing the frame count to $T_{a}$, $T_{b}$, and $T_{c}=T^{\prime}$ respectively.
Temporal Scene Clustering (TSC): The role of TSC is to perform visual content aware preliminary frame sampling. The established approach for preliminary frame selection is uniform sampling (limited to at most 200 frames). In contrast, TSC processes 900 to 1800 uniformly sampled frames to extract per-frame visual features using a lightweight deep neural network (ResNet-18) followed by a clustering procedure to identify $n$ non-overlapping frame sets. Within each of the n sets, we uniformly sample $\leq\tau$ frames obtaining a total of $T_{a}\leq\tau\times n$. Our iterative clustering procedure is outlined in Algorithm 1. It calculates pairwise distances between all frames accounting for intra-frame local information using the extracted per-frame features, followed by n iterative frame similarity based clustering operations. A single cluster could contain just one frame or significantly more based on frame feature similarities, leading to a non-uniform sampling of frames across the entire video. This allows more frames to be sampled from the information heavy temporal regions of videos.
Coarse Keyframe Detector (CKD): Unlike TSC in the prior stage, CKD reasons across both visual and language modalities (using the paired textual query, $q$) to further sub-sample $T_{a}$ into $T_{b}$ frames. CKD contains three elements: keyword generation strategy, dual-encoder image-text model, and similarity based confidence assignment algorithm. Keyword generation utilizes the given query, q, alongside hand-crafted templating operations or an LLM to select or generate suitable keywords. The dual-encoder image-text model uses a spatially aware contrastive language image pre-training (CLIP) network from (Ranasinghe et al., 2023). For confidence assignment, we construct an algorithm as outlined in Algorithm 2 which processes two lists, one of frames and one of keywords, and then calculates their pairwise likelihood of occurrence to assign each frame a confidence value (that reflects its usefulness to answer the query, $\mathbf{q}$). See  Appendix A.3 for more details.
@@
Our visual templating module partitions the $T_{b}$ frames into sets of 8 ordered by their confidence scores, arranges frame sets as grids to form a collage-style image, and annotates that image with visually identifiable tags corresponding to each frame. We further illustrate this process in Figure 3 (see Visual Templating column). Each of these visual templated images also contain a subset of keywords that correspond to their 8 images. These resulting visual templated images along with a prompt containing their associated keywords and instructions to select a frame subset based on valid association between keywords and images (see Appendix A.4 for details) are input to the VLM. The output of the VLM is used to select a subset of each 8 image group. These frames are collected as output of the FKD stage, overall resulting in $T_{c}$ frames.
The purpose of the initial visual templating module is to allow reasoning across a set of frames using the image-text VLM (which is trained to process a single image at time). This partitioning of the input $T_{b}$ frames is performed based on confidence scores from the prior stage and timestamps. The eight frames with top confidence scores are grouped into the first visual template, followed by the next eight and so forth. This ensures the VLM selects both high confidence concepts and low confidence concepts, accounting for biases and weaknesses in our CKD stage. After that, we temporally reorder some image sets with low confidence scores to cover keyframes distributed across long-range segments, while the sets with high confidence scores concentrate on keyframes in short-range segments. A total of 16 low-score frames are temporally reordered in this process. The algorithm is described in Algorithm 3 and the prompting technique is explained in Appendix A.4. Our intuition is that such a mechanism allows one to best utilize the complementary strengths of two different VLMs from CKD and FKD stages for better frame selection overall.
@@ 4 Experiments
In this section, we first discuss our experimental setup followed by quantitative evaluations comparing to existing baselines and ablations of our proposed components. We then present qualitative results for our method and outline some limitations of our approach.
Model EgoSchema NExT-QA IntentQA
Cap. Acc. (%) Cap. Acc. (%) Cap. Acc. (%)
Vamos (Wang et al., 2023) - 48.3 - - - -
IG-VLM (Kim et al., 2024) - 59.8 - 68.6 - 65.3
VideoLLaMA 2 (Cheng et al., 2024) - 53.3 - - - -
InternVideo2 (Wang et al., 2024c) - 60.2 - - - -
Tarsier (Wang et al., 2024a) - 61.7 - 79.2 - -
VIOLET (Fu et al., 2023) 5 19.9 - - - -
mPLUG-Owl (Ye et al., 2023) 5 31.1 - - - -
VideoAgent (Wang et al., 2024b) 8.4 54.1 8.2 71.3 - -
MVU (Ranasinghe et al., 2024) 16 37.6 16 55.2 - -
MoReVQA (Min et al., 2024) 30 51.7 16 69.2 - -
VFC (Momeni et al., 2023) - - 32 51.5 - -
SeViLA (Yu et al., 2024) 32 22.7 32 63.6 32 60.9
ProViQ (Choudhury et al., 2023) 60 57.1 60 64.6 - -
VideoTree (Wang et al., 2024e) 62.4 61.1 (56) 73.5 (56) 66.9
FrozenBiLM (Yang et al., 2022) 90 26.9 - - - -
LifelongMemory (Wang et al., 2024d) 90 62.1 - - - -
TraveLER (Shang et al., 2024) (101) 53.3 (65) 68.2 - -
LangRepo (Kahatapitiya et al., 2024) 180 41.2 90 60.9 90 59.1
LLoVi (Zhang et al., 2023) 180 50.3 90 67.7 90 64.0
LVNet (ours) 12 61.1 12 72.9 12 71.7
Table 2: Long Video Evalation: LVNet achieves state-of-the-art accuracies of 71.7%, 61.1%, and 72.9% on EgoSchema, NExT-QA, and IntentQA datasets respectively using just 12 frames compared to models using a similar number of captions. Models are ordered based on number of captions processed per video. Models with video-caption pretraining or utilizing significantly more captions than 12 frames used by LVNet are de-emphasized in grey or downplayed in light green to ensure fair comparison. Numbers in parentheses () indicate the maximum number of frames used. See Sec. A.2 in appendix for detailed results.
Metric Category VideoAgent VideoTree LVNet
Avg. Frames $\downarrow$ - 24.6 98.0 24.0
Acc.(%) $\uparrow$ Knowledge 52.2 60.7 63.0
Film & TV 42.5 52.5 45.0
Sports Comp. 42.7 48.6 48.0
Artistic Perf. 47.5 51.6 53.0
Life Record 44.7 49.5 45.0
Multilingual 36.6 40.0 53.0
@@
Average 46.4 53.1 52.4
Table 3: Evaluating on Very Long Videos. Comparison of LVNet (ours) with VideoAgent and VideoTree on the long split of VideoMME. LVNet uses the fewest frames while achieving the highest accuracy in three out of six categories and ranking second in overall performance, slightly below the best score. In the table, bold indicates the best performance, while underlined represents the second-best performance.

### v6

@@ Abstract
Long-form videos that span across wide temporal intervals are highly information redundant and contain multiple distinct events or entities that are often loosely related. Therefore, when performing long-form video question answering (LVQA), all information necessary to generate a correct response can often be contained within a small subset of frames. Recent literature leverage large language models (LLMs) in LVQA benchmarks, achieving exceptional performance, while relying on vision language models (VLMs) to convert all visual content within videos into natural language. Such VLMs often independently caption a large number of frames uniformly sampled from long videos, which is not efficient and can mostly be redundant. Motivated by this inefficiency, we propose LVNet, a modular and training-free framework featuring a novel Hierarchical Keyframe Selector (HKS) that efficiently selects a minimal set of informative frames tailored to each question. LVNet’s modularity allows easy integration with existing approaches for more efficient LVQA. We achieve state-of-the-art performance among similarly configured models across four benchmark LVQA datasets: EgoSchema, NExT-QA, IntentQA, VideoMME. The code can be found at https://github.com/jongwoopark7978/LVNet
@@ 1 Introduction
Video understanding is a long-standing vision problem (Aggarwal and Ryoo, 2011) with numerous real-world applications. It has been traditionally studied even before the era of differentiable representation learning, with hierarchical approaches focusing on longer videos (Allen and Ferguson, 1994; Ivanov and Bobick, 2000; Shi et al., 2004; Hongeng et al., 2004; Ryoo and Aggarwal, 2006). Today, video understanding research involving the language modality is particularly popular, with tasks such as video question answering (QnA) that involve generating human-style conversations using large language models (LLMs) (Tapaswi et al., 2016; Zeng et al., 2017; Xu et al., 2017).
Figure 1: High Accuracy with Low Compute: LVNet achieves state-of-the-art performance on EgoSchema (subset) while processing only a fraction of frame captions (below 12 per video) with the expensive LLM. More detailed analysis presented in Section˜$4.$
Motivated by successful image-language models (Liu et al., 2023; Dai et al., 2023), several works build video LLMs (Yu et al., 2023; Papalampidi et al., 2023; Maaz et al., 2023; Wang et al., 2024b) that directly process video frames in one stage to perform QnA. However, for long videos (containing 1000s of frames) these models require processing large visual token sequences with the LLM, making inference computationally expensive or even infeasible (see Table˜5). An alternate line of works (Zhang et al., 2023; Wang et al., 2023; Ranasinghe et al., 2024) follow multi-stage pipelines that first extract frame-level information followed by temporal modeling with an LLM. While these multi-stage works similarly encounter compute bottlenecks with increased frame processing for longer videos, it is possible to feed the LLM with descriptors of non-uniformly sampled frames Kahatapitiya et al. (2024); Wang et al. (2024d, f), motivating our exploration into keyframe selection, i.e. identifying a minimal set of frames most useful for correctly answering a given video-question pair.
@@
Therein, we propose LVNet, a framework containing a novel Hierarchical Keyframe Selector (HKS) that performs efficient key-frame selection followed by VLM for caption generation and LLM for answer generation as illustrated in Figure˜2. Aligned with prior work (Zhang et al., 2023; Wang et al., 2024f, d), per-frame captions are processed with a powerful LLM to generate correct answers for a given video-question pair. As shown in Fig. 1, LVNet achieves strong performance efficiently, processing only a fraction of frame descriptors (captions) with the LLM. Compared to feeding all captions or all frames directly to a powerful model (e.g. GPT-4o), our LVNet performs inference at a fraction of the cost (see Table˜6).
We summarize our key contributions as follows:
@@
    Efficient Frame Selection: Hierarchical Keyframe Selector (HKS) efficiently extracts keyframes from sequences up to 1800 frames (hour long videos) with a filter rate over 98%.
@@
    Video Training Free: Our framework requires no video level training and simply uses existing off-the-shelf modules.
@@
    Versatile Framework: Existing methods can easily be integrated with LVNet to further boost their performance (details in Appendix˜A.7).
@@
Proposed LVNet achieves state-of-the-art results (at common inference compute) on multiple long-form video question answering benchmarks demonstrating strong performance and generality.
@@ 2 Related Work
Video Question Answering: Visual question answering (VQA) involves generating open-ended textual content conditioned on an image and natural language query (Agrawal et al., 2015). Its video variant, Video-VQA (Yu et al., 2019a) replaces images with videos. Multiple early datasets focus on querying objects or events based on referential and spatial relations (Xu et al., 2017; Zeng et al., 2017; Yu et al., 2019a). Later tasks require explicit temporal understanding of sequential events (Lei et al., 2018, 2020; Yu et al., 2019b). More recent datasets focus on longer videos containing multiple actions and scenes spread over wide time intervals (termed long-form videos) (Xiao et al., 2021; Li et al., 2022). Referred to as long-form video question answering (LVQA), these benchmarks are constructed to specifically test strong causal and temporal reasoning (Xiao et al., 2021) over long temporal windows (Mangalam et al., 2023). Some works tackling such video VQA tasks leverage graph networks to model cross object / event relations (Hosseini et al., 2022; Xiao et al., 2022a, b). A more recent line of works integrate LLMs to tackle this task (Zhang et al., 2023; Wang et al., 2023; Kahatapitiya et al., 2024; Wang et al., 2024d; Ranasinghe et al., 2024; Wang et al., 2024f; Fan et al., 2024) utilizing the strong reasoning skills of LLMs. A common aspect is the use of a vision language model (VLM) to convert frame level visual information into natural language. This in turn is input to the LLM which makes a final prediction.
@@
Video Training Free
Table 1: LVNet exhibits unique features compared to prior work VideoAgent (VA) Fan et al. (2024), Traveler (Tr.) Shang et al. (2024), VideoTree (VT) Wang et al. (2024g), VideoChat-T (VC) Zeng et al. (2024), and Frame-Voyager (FV) Yu et al. (2024b) . See Appendix˜A.8 for details.
Frame Selection in Videos: The task of frame selection in videos has been long explored in video (Davis and Bobick, 1997; Zhao et al., 2017) with more recent works focused directly on long-form video question answering (Buch et al., 2022; Wang et al., 2024g; Fan et al., 2024; Zeng et al., 2024; Yu et al., 2024b). Most similar to our work is Wang et al. (2024d) which employs an LLM based strategy for video frame selection. However, our LVNet differs with several unique features as summarized in Table 1.
@@ 3 Method
Our proposed LVNet comprises of three components: a Hierarchical Keyframe Selector (HKS), a Vision Language Model (VLM), and a Large Language Model (LLM) as illustrated in Figure˜2. The HKS, an efficient, hierarchical keyframe selector, is the core contribution of our work. First, the model processes 900 uniformly sampled frames and clusters them into distinct scenes Next, it extracts keywords from a given natural language query via LLM and selects the frames most relevant to those keywords. Finally, the selected frames are described in natural language by a more powerful and computationally intensive VLM. Finally, an LLM processes the language descriptions of the selected frames to answer a given query.
@@ 3.1 Background
Recent approaches utilizing LLMs for long video question answering (LVQA) (Zhang et al., 2023; Wang et al., 2023; Kahatapitiya et al., 2024; Ranasinghe et al., 2024; Wang et al., 2024d) can be viewed as a composition of three sequential stages: a) frame selection, b) VLM based frame captioning, and c) LLM based answer generation. Note that the complexity of each stage varies across methods given their focus on different aspects of the LVQA task (e.g. frame selection in some is simply uniform sampling). In our work, we also follow this structure, but we focus on improving the frame selection stage. Under such a framework, our proposed HKS can serve as plug-in modules to replace the frame selection stage and the later two stages are similar to these prior works.
@@ 3.2 Architecture
Our LVNet processes a given video-query ($\mathbf{x},\mathbf{q}$) pair to output a response, $\mathbf{\hat{r}}$. The HKS module initially processes this video-query pair, selects $T^{\prime}$ keyframes, and outputs a deterministically sub-sampled video $\mathbf{x^{\prime}}\in\mathbb{R}^{T^{\prime}\times C\times H\times W}$. Each of these $T^{\prime}$ frames is then passed through the captioning stage of our VLM to generate a set of natural language descriptions, $D=\{d_{1},d_{2},...d_{T^{\prime}}\}$ where $d_{i}$ describes the frame $\mathbf{x^{\prime}}[i]$. Finally, the LLM processes all descriptions $D$ and the query $\mathbf{q}$ to generate response $\mathbf{\hat{r}}$. We illustrate this overall architecture in Figure˜2.
@@ 3.3 Hierarchical Keyframe Selector
We now describe our proposed Hierarchical Keyframe Selector (HKS) module. As illustrated in Figure˜2, our proposed HKS comprises of three sequential submodules, each reducing the frame count to $T_{a}$, $T_{b}$, and $T_{c}=T^{\prime}$ respectively.
Temporal Scene Clustering (TSC): The role of TSC is to perform visual content aware preliminary frame sampling. The established approach for preliminary frame selection is uniform sampling (limited to at most 200 frames). In contrast, TSC processes 900 to 1800 uniformly sampled frames to extract per-frame visual features using a lightweight deep neural network (ResNet-18) followed by a clustering procedure to identify $n$ non-overlapping frame sets. Within each of the n sets, we uniformly sample $\leq\tau$ frames obtaining a total of $T_{a}\leq\tau\times n$. Our iterative clustering procedure is outlined in Algorithm˜1. It calculates pairwise distances between all frames accounting for intra-frame local information using the extracted per-frame features, followed by n iterative frame similarity based clustering operations. A single cluster could contain just one frame or significantly more based on frame feature similarities, leading to a non-uniform sampling of frames across the entire video. This allows more frames to be sampled from the information heavy temporal regions of videos.
Coarse Keyframe Detector (CKD): Unlike TSC in the prior stage, CKD reasons across both visual and language modalities (using the paired textual query, $q$) to further sub-sample $T_{a}$ into $T_{b}$ frames. CKD contains three elements: keyword generation strategy, dual-encoder image-text model, and similarity based confidence assignment algorithm. Keyword generation utilizes the given query, q, alongside hand-crafted templating operations or an LLM to select or generate suitable keywords. The dual-encoder image-text model uses a spatially aware contrastive language image pre-training (CLIP) network from (Ranasinghe et al., 2023). For confidence assignment, we construct an algorithm as outlined in Algorithm˜2 which processes two lists, one of frames and one of keywords, and then calculates their pairwise likelihood of occurrence to assign each frame a confidence value (that reflects its usefulness to answer the query, $\mathbf{q}$). See  Appendix˜A.5 for more details.
@@
Our visual templating module partitions the $T_{b}$ frames into sets of 8 ordered by their confidence scores, arranges frame sets as grids to form a collage-style image, and annotates that image with visually identifiable tags corresponding to each frame. We further illustrate this process in Figure˜3 (see Visual Templating column). Each of these visual templated images also contain a subset of keywords that correspond to their 8 images. These resulting visual templated images along with a prompt containing their associated keywords and instructions to select a frame subset based on valid association between keywords and images (see Appendix˜A.6 for details) are input to the VLM. The output of the VLM is used to select a subset of each 8 image group. These frames are collected as output of the FKD stage, overall resulting in $T_{c}$ frames.
The purpose of the initial visual templating module is to allow reasoning across a set of frames using the image-text VLM (which is trained to process a single image at time). This partitioning of the input $T_{b}$ frames is performed based on confidence scores from the prior stage and timestamps. The eight frames with top confidence scores are grouped into the first visual template, followed by the next eight and so forth. This ensures the VLM selects both high confidence concepts and low confidence concepts, accounting for biases and weaknesses in our CKD stage. After that, we temporally reorder some image sets with low confidence scores to cover keyframes distributed across long-range segments, while the sets with high confidence scores concentrate on keyframes in short-range segments. A total of 16 low-score frames are temporally reordered in this process. The algorithm is described in Algorithm˜3 and the prompting technique is explained in Appendix˜A.6. Our intuition is that such a mechanism allows one to best utilize the complementary strengths of two different VLMs from CKD and FKD stages for better frame selection overall.
@@ 4 Experiments
In this section, we first discuss our experimental setup followed by quantitative evaluations comparing to existing baselines and ablations of our proposed components. We then discuss the results on efficiency and scalability, and finally outline limitations.
Model EgoSchema NExT-QA IntentQA VT Free
@@
VideoLLaMA 2 (Cheng et al., 2024) - 53.3 - - - - no
InternVideo2 (Wang et al., 2024e) - 60.2 - - - - no
Tarsier (Wang et al., 2024a) - 61.7 - 79.2 - - no
Vamos (Wang et al., 2023) - 48.3 - - - - yes
IG-VLM (Kim et al., 2024) - 59.8 - 68.6 - 65.3 yes
VIOLET (Fu et al., 2023) 5 19.9 - - - - yes
mPLUG-Owl (Ye et al., 2023) 5 31.1 - - - - yes
VideoAgent (Wang et al., 2024d) 8.4 54.1 8.2 71.3 - - yes
MVU (Ranasinghe et al., 2024) 16 37.6 16 55.2 - - yes
MoReVQA (Min et al., 2024) 30 51.7 16 69.2 - - yes
VFC (Momeni et al., 2023) - - 32 51.5 - - yes
ProViQ (Choudhury et al., 2023) 50 57.1 50 64.6 - - yes
VideoTree (Wang et al., 2024g) 62.4 61.1 (56) 73.5 (56) 66.9 yes
FrozenBiLM (Yang et al., 2022) 90 26.9 - - - - yes
LifelongMemory (Wang et al., 2024f) 90 62.1 - - - - yes
TraveLER (Shang et al., 2024) (101) 53.3 (65) 68.2 - - yes
LangRepo (Kahatapitiya et al., 2024) 180 41.2 90 60.9 90 59.1 yes
LLoVi (Zhang et al., 2023) 180 50.3 90 67.7 90 64.0 yes
LVNet (ours) 12 61.1 12 72.9 12 71.7 yes
Table 2: Long Video Evalation: LVNet achieves state-of-the-art accuracies of 61.1%, 72.9%, and 71.7% on EgoSchema, NExT-QA, and IntentQA datasets respectively using just 12 frames compared to models using a similar number of captions. Models are ordered based on number of captions processed per video. Models with video-level training (VT) or utilizing significantly more captions than 12 frames used by LVNet are de-emphasized in grey or downplayed in light green to ensure fair comparison. Numbers in parentheses () indicate the maximum number of frames used. See Sec. A.2 in appendix for detailed results.
[…]
VideoChat-T 7B/OS no N/A 43.8
Frame-Voyager 34B/OS no N/A 51.2
LVNet (DS-V3) 37B/OS yes 24 53.1
VideoAgent+GPT-4o <1.8T/$P$ yes 24 51.3
VideoTree+GPT-4o <1.8T/$P$ yes 24 52.3
LVNet (GPT-4o) <1.8T/$P$ yes 24 53.9
Table 3: Comparison to Keyframe Selection Methods on VideoMME: We compare LVNet with both single-stage methods that rely on video-level training (VideoChat-T, Frame-Voyager) and two-stage methods (VideoAgent, VideoTree). For the two-stage comparison, we evaluate VideoTree and VideoAgent with GPT-4o as both the captioner and LLM under the same 24-frame caption budget for a direct head-to-head comparison. Legend: DS-V3=DeepSeek-V3, OS=open-source, PP=proprietary, VT Free=no video-level training, TS Cap.=two-stage caption numbers.. Notably, LVNet outperforms both single-stage and two-stage methods without requiring video-level training.